Subtract the polynomials.
(8n^3-n^7-12)-(7n^3-n^7-8)
Can some one break this down for me step by step so I can see how it is done. Thank you for your time.

Answer by tommyt3rd(5050)

We convert to addition by reversing the operation on every term of the subtrahend (the polynomial being "taken away"). Then we combine like terms:

(8n^3-n^7-12)-(7n^3-n^7-8)=
8n^3-n^7-12-7n^3+n^7+8=
n^3-4

:)